What are the vertices of 81x^2+9y^2=49
Answer by stanbon(75887)

What are the vertices of 81x^2+9y^2=49
------
[x^2/(49/81)] + y^2/[49/9] = 1
-----------
[x^2/(7/9)^2] + y^2/[(7/3)^2] = 1
-------
Center at (0,0)
Vertices on major axis: (0,-7/3) ; (0,+7/3)
Vertices of minor axis: (-7/9,0) ; (7/9,0)
=============================
Cheers,
Stan H.
====================